# Distilling Diffusion Models to Efficient 3D LiDAR Scene Completion

### v1

@@ 4.3 Optimization procedure
We set $\lambda_{scene}=0.5$ and $\lambda_{point}=0.01$ unless otherwise specified. The implementation details are provided in Sec. A.3.
Model CD $\downarrow$ JSD $\downarrow$ Times (s) $\downarrow$
LMSCNet [26] 0.641 0.431 0.40
LODE [14] 1.029 0.451 1.76
MID [34] 0.503 0.470 6.42
PVD [51] 1.256 0.498 -
LiDiff [24] 0.434 0.444 30.38
LiDiff (Refined) [24] 0.375 0.416 30.55
ScoreLiDAR 0.406 0.425 5.16
ScoreLiDAR (Refined) 0.342 0.399 5.37
[…]
Model CD $\downarrow$ JSD $\downarrow$ Times (s) $\downarrow$
LMSCNet [26] 0.979 0.496 -
LODE [14] 1.565 0.483 -
MID [34] 0.637 0.476 -
LiDiff [24] 0.564 0.459 29.18
LiDiff (Refined) [24] 0.517 0.446 29.43
ScoreLiDAR 0.472 0.444 4.98
ScoreLiDAR (Refined) 0.452 0.437 5.14
@@
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
ScoreLiDAR (Refined) 0.342 0.399 0.452 0.437
w/o Structural Loss 0.419 0.430 0.549 0.445
@@
Model CD $\downarrow$ JSD $\downarrow$ Time (s) $\downarrow$
LiDiff (50 steps) [24] 0.434 0.444 30.38
LiDiff (50 steps Refined) [24] 0.375 0.416 30.55
LiDiff (8 steps) [24] 0.447 0.432 5.69
LiDiff (8 steps Refined) [24] 0.411 0.406 5.92
ScoreLiDAR (8 Steps Refined) 0.342 0.399 5.37
ScoreLiDAR (4 Steps Refined) 0.326 0.386 3.23
ScoreLiDAR (2 Steps Refined) 0.403 0.379 1.85
ScoreLiDAR (1 Steps Refined) 0.548 0.384 1.10
@@ 5.1 Scene completion
We validate ScoreLiDAR on SemanticKITTI [1] and KITTI-360 [17] datasets. The existing SOTA LiDAR scene completion model LiDiff [24] is chosen as the teacher model. The student model shares the network architecture with the teacher model and is initialized by the teacher model. Moreover, we also use the refinement network in LiDiff [24] to refine the completed scene generated by the student model. We calculate the Chamfer Distance (CD) and the Jensen-Shannon Divergence (JSD) to evaluate the similarity between the completed scene and the ground truth. The smaller the value of CD and JSD, the closer the completed scene is to the ground truth.
Tab. 1 shows the quantitative results on SemanticKITTI. ScoreLiDAR achieves the optimal performance compared to the existing models on both metrics. Compared to the SOTA method LiDiff [24] with refinement, which takes 30.55 seconds to complete a scene, our proposed ScoreLiDAR completes a scene in just 5.47 seconds (fivefold speedup) yet with $8\%$ improvement in CD and $4\%$ in JSD. Although LMSCNet [26] and LODE [14] have faster completion speeds, their completion quality is significantly lower. Notably, the performance of ScoreLiDAR outperforms the teacher model LiDiff [24]. This is because ScoreLiDAR introduces a structural loss with scene-wise term and point-wise term, enabling the student model to effectively capture geometric structure information within LiDAR point cloud data during training. Results on KITTI-360 are shown in Tab. 2. ScoreLiDAR also achieves optimal performance compared to existing models. ScoreLiDAR also boasts a fivefold speedup with $12\%$ improvement in CD and $2\%$ in JSD compared to LiDiff [24].
@@ 5.2 Ablation study
In this part, we conduct the ablation study to verify the effectiveness of the structural loss in the training of the proposed ScoreLiDAR. We compared the scene completion performances of the proposed ScoreLiDAR with a variant that does not incorporate structural loss. The results are shown in Tab. 3. The results show that the variant without structural loss exhibits lower performance in scene completion on both datasets. However, after considering the structural loss, the performance of ScoreLiDAR improves significantly, which achieves better performance on both metrics. This supports our discussion in Sec. 4.3, incorporating structural loss enables the student model to capture the geometric structure feature of 3D point clouds, thereby facilitating the effective distillation of the student model.
Furthermore, we compared the scene completion performance of ScoreLiDAR with different sampling steps, and the results are shown in Tab. 4. It can be observed that as the sampling steps decrease from 8 to 1, the time required for ScoreLiDAR to complete a scene also decreases, with single-step sampling allowing a scene to be completed in only 1.1 seconds. With 8-step and 4-step sampling, ScoreLiDAR performs better on both metrics than LiDiff. Both metrics decay at 2-step and 1-step sampling, but in JSD ours still performs better than LiDiff. In summary, although the quality of scene completion decreases as the sampling steps are reduced, it still maintains performance comparable to or better than the existing model, achieving better performance and speed trade-off as in Fig. 2.
@@
5.3 Qualitative analysis
Fig. 4 shows the completed scenes by our proposed ScoreLiDAR and LiDiff [24] on KITTI-360. ScoreLiDAR achieves completion results with higher quality and greater fidelity. We can see that LiDiff [24] nearly fails at 8 steps. Although LiDiff [24] achieves decent completion at 50 steps, there are some missing areas on the lower and right sides. In contrast, ScoreLiDAR reaches optimal completion including clearer objects such as cars, traffic cones, and other scene elements with only 8 steps, which is closer to the ground truth, especially after refinement.
@@ 5.3 Qualitative analysis
We also conduct a user study to evaluate the performance of ScoreLiDAR. We used ScoreLiDAR and LiDiff [24] to complete scenes based on the same input scans and asked users to choose the scene they believed was closer to the ground truth. ScoreLiDAR received a $65\%$ user preference over LiDiff [24]. This indicates that the detail and fidelity of the scenes completed by ScoreLiDAR more closely resemble the ground truth for most users. The details of the user study are shown in Sec. D.4.
@@ 6 Conclusion
This paper proposes ScoreLiDAR, a novel distillation method tailored for 3D LiDAR scene completion diffusion models. The distilled model enables efficient LiDAR scene completion. By introducing the structural loss with scene-wise term and point-wise term, ScoreLiDAR trains the student model to effectively capture the holistic structure and the relative configuration of key points and achieve efficient and high-quality scene completion.
@@
While ScoreLiDAR achieves efficient, high-quality LiDAR scene completion, its performance is constrained by the teacher model. As the performance of the teacher model improves, so does the capability of the student model. Moreover, the ability to perform semantic scene completion is also determined by the teacher model. If the teacher model can complete the semantic scene, the distilled student model will also be able to follow the teacher. Thus, further exploration is required to find a more effective method to improve the training process [45, 44] of ScoreLiDAR and avoid the limitations of the teacher model, achieving more efficient semantic LiDAR scene completion.
@@ A.1 Dataset setup
Figure 6: Comparison of scene completion details between ScoreLiDAR and the SOTA model LiDiff [24]. The magnified images are enlarged views of the regions corresponding to the boxes in the completed scene images. Compared to LiDiff, ScoreLiDAR better completes the details of vehicles, making it closer to the ground truth scene.
@@ A.3 Implementation details
We choose the pre-trained LiDiff [24] model as the teacher model $\boldsymbol{\epsilon_{\theta}}$, the student model $G_{stu}$ and the auxiliary diffusion model $\boldsymbol{\epsilon}_{\phi}$ shares the same network architecture as the teacher model and are initialized by the teacher model. The ScoreLiDAR is trained on SemanticKITTI dataset. The pre-trained diffusion model is provided by the official release of LiDiff [24].
For optimization, we use the Stochastic Gradient Descent (SGD) optimizer with the default parameters. The learning rate is set to $53e-$ 5 and the batch size is set to $1$. The training ratio between the student model and the auxiliary diffusion model is maintained at $1:1$. To reduce computational costs, when calculating the point-wise loss, we first randomly select $\frac{1}{10}$ of the points from the ground truth scene. Then, following the proposed method, we select the top $\frac{1}{3}$ points with the highest curvature from these points as the key points to calculate the distance matrix. That is, the final number of key points is $\frac{1}{30}$ of the total number of points in the ground truth scene. When calculating the $K$-nearest neighbours, we set $K=180$. The weights of scene-wise loss $\lambda_{scene}$ and the point-wise loss $\lambda_{point}$ are set to $0.5$ and $0.01$, respectively. ScoreLiDAR requires only 50 iterations to achieve convergence, taking approximately 10 minutes on a single A40 GPU, which is highly efficient. Our code will be publicly available soon.
@@ Appendix B Discussion
Firstly, we discuss the significance of this study. For autonomous vehicles, accurately recognizing and perceiving their surrounding environment during operation is critical [16, 15]. This is particularly important for identifying objects that may affect the vehicle’s movement, such as other vehicles, pedestrians, traffic cones, and signposts [11]. Accurate and efficient recognition of these objects is essential for the safe operation of autonomous vehicles. However, the scan data obtained by onboard LiDAR as shown in Fig. 6 (a) is sparse [34, 14], and it is difficult to identify key objects such as vehicles from the magnified regions of the sparse scan. However, from the ground truth scene in Fig. 6 (b), it is evident that these regions contain moving vehicles. Autonomous vehicles cannot obtain sufficient information about the driving environment from these sparse LiDAR scans [25, 13]. Therefore, it is necessary to use appropriate methods to complete the sparse LiDAR scans.
@@
LiDiff [24] uses DDPM [9] models to complete 3D LiDAR scenes, achieving impressive results. However, due to the inherent characteristics of diffusion models, LiDiff [24] requires approximately 30 seconds to complete a single scene, limiting its applicability in autonomous vehicles. In contrast, the proposed ScoreLiDAR can complete a scene in almost 5 seconds, more than 5 times faster than LiDiff [24], while achieving higher completion quality. As shown in Fig. 6 (c), in the magnified region, although LiDiff [24] completes some vehicle shapes, they are not fully reconstructed. For example, in the left region, there should be two vehicles, but the scene completed by LiDiff only contains one. In contrast, the scene completed by ScoreLiDAR features clearer and more complete vehicle structures (Fig. 6 (d)), making it closer to the ground truth. Thus, with the scenes completed by the proposed ScoreLiDAR, autonomous vehicles can more easily recognize critical objects in their driving environment, enabling safer and more effective navigation.
Moreover, ScoreLiDAR allows for different sampling steps during distillation to achieve varying completion speeds. However, due to the limitations of the teacher model, ScoreLiDAR’s completion quality decreases at smaller sampling steps. With further improvements or by replacing the teacher model for better distillation, ScoreLiDAR could complete a scene within one second, achieving real-time scene completion. This is crucial for autonomous vehicles to scan and recognize their driving environment effectively. Therefore, future work will focus on exploring this aspect.

### v2

@@ 4.3 Optimization procedure
We set $\lambda_{scene}=0.5$ and $\lambda_{point}=0.01$ defaultly. The implementation details are provided in Sec S1.
Model CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ Times (s) $\downarrow$
LMSCNet [30] 0.641 0.431 - 0.40
LODE [16] 1.029 0.451 - 1.76
MID [38] 0.503 0.470 - 6.42
PVD [55] 1.256 0.498 - 262.54
LiDiff [28] 0.434 0.444 22.15 30.38
LiDiff (Refined) [28] 0.375 0.416 23.16 30.55
ScoreLiDAR 0.406 0.425 23.14 5.16
ScoreLiDAR (Refined) 0.342 0.399 23.26 5.37
Table 1: The completion performance on the SemanticKITTI dataset. Colors denote the 1st, 2nd, and 3rd best-performing model. The sampling time is estimated based on the official code and the provided checkpoints.
Model CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ Times (s) $\downarrow$
LMSCNet [30] 0.979 0.496 - 0.38
LODE [16] 1.565 0.483 - 1.64
MID [38] 0.637 0.476 - 6.23
LiDiff [28] 0.564 0.459 21.98 29.18
LiDiff (Refined) [28] 0.517 0.446 22.96 29.43
ScoreLiDAR 0.472 0.444 22.80 4.98
ScoreLiDAR (Refined) 0.452 0.437 23.02 5.14
@@
CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$
ScoreLiDAR (Refined) 0.342 0.399 23.26 0.452 0.437 23.02
w/o Structural Loss (Refined) 0.419 0.430 24.61 0.549 0.445 24.56
@@
Model CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ Time (s) $\downarrow$
LiDiff (50 steps) [28] 0.434 0.444 22.15 30.38
LiDiff (50 steps Refined) [28] 0.375 0.416 23.16 30.55
LiDiff (8 steps) [28] 0.447 0.432 24.90 5.69
LiDiff (8 steps Refined) [28] 0.411 0.406 25.74 5.92
ScoreLiDAR (8 Steps Refined) 0.342 0.399 23.14 5.37
ScoreLiDAR (4 Steps Refined) 0.326 0.386 23.98 3.23
ScoreLiDAR (2 Steps Refined) 0.403 0.379 - 1.85
ScoreLiDAR (1 Steps Refined) 0.548 0.384 - 1.10
@@ 5.1 Scene completion
We validate ScoreLiDAR on SemanticKITTI [1] and KITTI-360 [20] datasets. The existing SOTA LiDAR scene completion model LiDiff [28] is chosen as the teacher model. The student model shares the network architecture with the teacher model and is initialized by the teacher model. Moreover, we also use the refinement network in LiDiff [28] to refine the completed scene generated by the student model. We calculate the Chamfer Distance (CD) [2], the Jensen-Shannon Divergence (JSD) [24] and the Earth Mover’s Distance (EMD) [8] to evaluate the similarity between the completed scene and the ground truth. The smaller the value of these metrics, the closer the completed scene is to the ground truth.
Tab. 1 shows that ScoreLiDAR achieves the optimal performance compared to the existing models in most cases on the SemanticeKITTI dataset. Compared to the SOTA method LiDiff [28] with refinement, which takes 30.55 seconds to complete a scene, ScoreLiDAR completes a scene in just 5.47 seconds (fivefold speedup) yet with $8\%$ improvement in CD, $4\%$ in JSD, and comparable results in EMD. The reason for the increase in EMD after refinement is the mismatch in the points numbers between $\mathcal{G}^{0}$ and $\mathcal{G}$, leading to a higher matching cost. The refinement process optimizes local details but alters the global point cloud distribution, extending the overall transfer path. Although LMSCNet [30] and LODE [16] have faster completion speeds, their completion quality is significantly lower. The performance of ScoreLiDAR outperforms the teacher model LiDiff [28]. This is because ScoreLiDAR introduces a structural loss with scene-wise term and point-wise term, enabling the student model to effectively capture geometric structure information within LiDAR point cloud data during training. Results on KITTI-360 are shown in Tab. 2. ScoreLiDAR also achieves optimal performance in most cases and boasts a fivefold speedup with $12\%$ improvement in CD and $2\%$ in JSD compared to LiDiff [28].
@@ 5.2 Ablation study
In this part, we conduct the ablation study to verify the effectiveness of the structural loss in the training of the proposed ScoreLiDAR. We compared the scene completion performances of the proposed ScoreLiDAR with a variant that does not incorporate structural loss. The results are shown in Tab. 3. The results show that the variant without structural loss exhibits lower performance in scene completion on both datasets. However, after considering the structural loss, the performance of ScoreLiDAR improves significantly, which achieves better performance on all metrics. This supports our discussion in Sec. 4.3, incorporating structural loss enables the student model to capture the geometric structure feature of 3D point clouds, thereby facilitating the effective distillation of the student model.
Furthermore, we compared the scene completion performance of ScoreLiDAR with different sampling steps, and the results are shown in Tab. 4. It can be observed that as the sampling steps decrease from 8 to 1, the time required for ScoreLiDAR to complete a scene also decreases, with single-step sampling allowing a scene to be completed in only 1.1 seconds. With 8-step and 4-step sampling, ScoreLiDAR performs better than LiDiff. All metrics decay at 2-step and 1-step sampling, but in JSD ours still performs better than LiDiff. In summary, although the quality of scene completion decreases as the sampling steps are reduced, it still maintains performance comparable to or better than the existing model, achieving better performance and speed trade-off as in Fig. 2.
In addition, we compared the ablation results about the terms of structural loss, different keypoint selection methods, varying numbers of keypoints, and different values of $\lambda_{scene}$ and $\lambda_{point}$ and so on, which are presented in Sec. S4.
5.3 Qualitative analysis
Fig. 4 shows the completed scenes by our proposed ScoreLiDAR and LiDiff [28] on KITTI-360. ScoreLiDAR achieves completion results with higher quality and greater fidelity in less time compared to LiDiff. For example, as shown in Fig. 4(c), in the left region, there should be two vehicles as in Fig. 4(b), but the scene completed by LiDiff only contains one. In contrast, the scene completed by ScoreLiDAR with only 8 steps in Fig. 4(d) completes two vehicles and has clearer and more complete vehicle structures, making it closer to the ground truth.
@@ 5.3 Qualitative analysis
We also conduct a user study to evaluate the performance of ScoreLiDAR. We used ScoreLiDAR and LiDiff [28] to complete scenes based on the same input scans and asked users to choose the scene they believed was closer to the ground truth. ScoreLiDAR received a $65\%$ user preference over LiDiff [28]. This indicates that the detail and fidelity of the scenes completed by ScoreLiDAR more closely resemble the ground truth for most users. The details of the user study are shown in SM.
@@ 6 Conclusion
This paper proposes ScoreLiDAR, a novel distillation method tailored for 3D LiDAR scene completion diffusion models based on a bidirectional gradient guidance mechanism. By introducing the structural loss with scene-wise term and point-wise term, ScoreLiDAR trains the student model to effectively capture the holistic structure and the relative configuration of key points and achieve efficient and high-quality scene completion.
@@
While ScoreLiDAR achieves efficient, high-quality LiDAR scene completion, its performance is constrained by the teacher model. As the performance of the teacher model improves, so does the capability of the student model. Although we conducted a preliminary experiment on the semantic scene completion in the Supplementary Material, a more comprehensive exploration is still required. Thus, further exploration is required to find a more effective method to improve the training process of ScoreLiDAR and avoid the limitations of the teacher model, achieving more efficient semantic LiDAR scene completion.
Acknowledgement
This paper is supported by Provincial Key Research and Development Plan of Zhejiang Province under No. 2024C01250(SD2), National Natural Science Foundation of China (Grant No. 62006208) and Dream Set Off - Kunpeng &\&& Ascend Seed Program.
@@ A.3 Implementation details
We choose the pre-trained LiDiff [28] model as the teacher model $\boldsymbol{\epsilon_{\theta}}$, the student model $G_{stu}$ and the auxiliary diffusion model $\boldsymbol{\epsilon}_{\phi}$ shares the same network architecture as the teacher model and are initialized by the teacher model. The ScoreLiDAR is trained on SemanticKITTI dataset. The pre-trained diffusion model is provided by the official release of LiDiff [28]. For fair comparison, we follow LiDiff [28]’s training strategy. ScoreLiDAR is trained on sequences 00–07 and 09–10 of SemanticKITTI (does not train on KITTI-360), and evaluated on sequence 08 of SemanticKITTI and sequence 00 of KITTI-360.
For optimization, we use the Stochastic Gradient Descent (SGD) optimizer with the default parameters. The learning rate is set to $53e-$ 5 and the batch size is set to $1$. The training ratio between the student model and the auxiliary diffusion model is maintained at $1:1$. To reduce computational costs, when calculating the point-wise loss, we first randomly select $\frac{1}{10}$ of the points from the ground truth scene. Then, following the proposed method, we select the top $\frac{1}{3}$ points with the highest curvature from these points as the key points to calculate the distance matrix. That is, the final number of key points is $\frac{1}{30}$ of the total number of points in the ground truth scene. When calculating the $K$-nearest neighbours, we set $K=180$. The weights of scene-wise loss $\lambda_{scene}$ and the point-wise loss $\lambda_{point}$ are set to $0.5$ and $0.01$, respectively. ScoreLiDAR requires only 50 iterations to achieve convergence, taking approximately 10 minutes on a single A40 GPU, which is highly efficient. Our model and code are publicly available on https://github.com/happyw1nd/ScoreLiDAR.
@@ Appendix B Discussion
B.1 The effectiveness of the distillation on improving the completion efficiency
In this part, we provide a detailed discussion about the efficiency of the proposed distillation method.
Why is it reasonable to initialize the student model and auxiliary diffusion model using the teacher model?
Firstly, such an initialization method is commonly used in existing methods [22, 54, 52, 51, 37]. Secondly, the pre-trained teacher model $\boldsymbol{\epsilon}_{\theta}$ contains the information about the training distribution, initializing the student model with the teacher model to perform distillation is essentially a fine-tuning process for the teacher model, which can accelerate the efficiency of the distillation. Third, although the student model has the same parameters and the network structure as the teacher model, its sampling distribution is different from that of the teacher model due to the different sampling steps. The teacher model $\boldsymbol{\epsilon}_{\theta}$ (LiDiff [28] in this case) conducts the multi-step sampling.
$\mathcal{G}^{t-1}=\frac{1}{\sqrt{\alpha^{t}}}\left(\mathcal{G}^{t}-\frac{1-\alpha^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)\right)+\sigma^{t}\boldsymbol{z}$ (21)
The teacher model $\boldsymbol{\epsilon}_{\theta}$ from LiDiff [28] conducts 50-steps sampling by repeating Eq. 21. The student model $G_{stu}$ conducts the single-step sampling. After $G_{stu}$ predicts the noise, a single-step denoising in Eq. 22 is performed to directly obtain the completed scene $\mathcal{G}^{0}$.
$\mathcal{G}^{0}=\frac{1}{\sqrt{\alpha^{t}}}\left(\mathcal{G}^{t}-\frac{1-\alpha^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)\right)$ (22)
Thus, the single-step sampling scene of the student model is different from the multi-step sampling scene of the teacher model.
Why is the distillation loss effective? Why does the student model get optimized?
Firstly, our distillation loss is different from the standard loss of the diffusion models such as DDPM [10]. The loss of DDPM is to directly predict the noise added to the training sample. Differently, in proposed ScoreLiDAR, the distillation loss utilizes the noise predicted by the student model $G_{stu}$ to perform a one-step sampling, resulting in a completed scene $\mathcal{G}^{0}$ different from the multi-step sampling of the teacher model. The completed scene $\mathcal{G}^{0}$ is then perturbed noise on a random timestep $t$ to obtain the noisy scene $\mathcal{G}^{t}$, and the difference between two score functions according to $\mathcal{G}^{t}$ is calculated to serve as the gradient for optimizing the student model, as shown in Eq. 23
[…]
Why is one-step sampling used during training, while few-step sampling is used during inference?
[…]
B.2 The differents of scene-wise loss and Chamfer Distance
The scene-wise loss has the following form
$\mathcal{L}_{scene}=\frac{1}{|\mathcal{G}^{0}|}\sum_{\boldsymbol{p}^{0}_{i}\in\mathcal{G}^{0}}\min_{\boldsymbol{p}\in\mathcal{G}}\|\boldsymbol{p}^{0}_{i}-\boldsymbol{p}\|^{2}$ (24)
The scene-wise loss is part of the Chamfer Distance. The reason for using only part of the Chamfer Distance (CD) is that the optimization objective of scene-wise loss is to ensure that the points in the generated scene $\mathcal{G}^{0}$ are as close as possible to the corresponding points in the ground truth. The unused term in CD matches each point in ground truth with its nearest point in generated scene $\mathcal{G}^{0}$, which may lead to points in $\mathcal{G}^{0}$ being pushed toward the average position of non-existent matching points in the ground truth. This effect is detrimental to scene completion.
B.3 Discussion on the significance of this study
@@
Firstly, we discuss the significance of this study. For autonomous vehicles, accurately recognizing and perceiving their surrounding environment during operation is critical [18, 17]. This is particularly important for identifying objects that may affect the vehicle’s movement, such as other vehicles, pedestrians, traffic cones, and signposts [12]. The accurate and efficient recognition of these objects is essential for the safe operation of autonomous vehicles. However, the scan data obtained by onboard LiDAR is sparse [38, 16], and it is difficult to identify key objects such as vehicles from the magnified regions of the sparse scan. Autonomous vehicles cannot obtain sufficient information about the driving environment from these sparse LiDAR scans [29, 14]. Therefore, it is necessary to use appropriate methods to complete the sparse LiDAR scans.
@@
LiDiff [28] uses DDPM [10] models to complete 3D LiDAR scenes, achieving impressive results. However, due to the inherent characteristics of diffusion models, LiDiff [28] requires approximately 30 seconds to complete a single scene, limiting its applicability in autonomous vehicles. In contrast, the proposed ScoreLiDAR can complete a scene in almost 5 seconds, more than 5 times faster than LiDiff [28], while achieving higher completion quality. Thus, with the scenes completed by the proposed ScoreLiDAR, autonomous vehicles can more easily recognize critical objects in their driving environment, enabling safer and more effective navigation.